# TRIM: Token Reduction and Inference Modeling for Cost-Effective Language Generation

### v3

@@ 7.1 Inferable terms set
In this study, we analyze the degradation of the metrics as the set of function words to be distilled and reconstructed increases. This will allow us to empirically decide a good choice of set $S$ for the second experiment. We applied Algorithm  1 using as the text corpus a random sample of 100,000 paragraphs from the Wikipedia dump, as a word set an extended list of 127 function words of the English language Jurafsky and Martin (2000), and RoBERTa as the language model.
@@
To determine the boundary for the subset of words that will be reconstructed a posteriori $S$, we followed an iterative procedure in which the T5 model is fine-tuned for the text reconstruction task using the top-N inferable terms from the sorted list, increasing the subset in 5 words, from size 5 up to 60 (12 levels). For this training we used a new random sample from the Wikipedia dump. Each training set consists of 100,000 (original, distilled) paragraph pairs where the distilled text is built by removing the words of the corresponding level. Additional 5,000 pair for testing were built with the same schema for each of the 12 levels. The performance is then evaluated at each of these distillation levels, analyzing the degradation of the results as the set size increases.
@@
The results per level can be seen in Figure 2, representing the values of the metrics against the percentage of saved tokens per level, and extrapolating trend lines. We can observe that from level 5 onward (around 18% of saved tokens), the percentage of tokens saved increases to a lesser extent while the metrics continue to degrade. Therefore, we set our boundary at level 5, and collect these 25 function words (set $S$) for the final evaluation.
@@ 7.2 Reconstruction model
Once we have segmented the set of inferable terms for this experimentation, we can focus on the development of the reconstruction model $M_{R}$ in charge of inferring the terms omitted from the generated text. For this purpose, we have taken a new random sample of 100,000 paragraphs from the Wikipedia dump, and we have removed the inferable terms of set $S$, thus forming the associated training corpus. The 4 language models selected have been fine-tuned in the text reconstruction task. The training information can be found in Table 4 of the Supplementary Material.
We tested each model on two test sets. First, on a new random sample of the Wikipedia dump to validate the training, and second, on the DKA dataset (Section 6). We evaluated their performance in a scenario closer to the real one where these models are used after the generation of a text omitting the inferable terms by an LLM. Both test datasets are composed of 9,000 paragraphs each. The results can be seen in Table 5 and 3, showing also the metrics of the distilled text compared to the original one.
Regarding the $\Theta$ metrics, we observe good results overall. The smaller language models Bart and T5 excel in precision, and the instruction-bigger language model Qwen2.5 in its 0.5B and 1.5B versions excel in recall. This means the smaller models tend to be more conservative but more precise, while on the other hand the bigger models reconstruct more terms from the original structure but with higher variability in their correctness. Nevertheless, the metrics are very equated, and do not reflect a high difference between them.
Regarding the rest of metrics, we can observe that the grammatical metrics (BLEU, METEOR & ROUGE) have good values across all models, the semantic metric is perfect with 3 decimals approximation, and the perplexity suffers only a small increase over the original one. All of these results are obtained with a significant percentage of saved tokens of 18.85% in the sample of Wikipedia and 20.58% for the answers of the LLM.
We can observe from the metrics for the distilled text how the grammatical metrics are substantially worse, and the perplexity much higher while the semantics metrics remain almost perfect. This shows that the elimination of the inferable sets breaks the grammatical structure and coherence but maintains almost in its fullness the original semantic information.
Focusing on the differences between test sets (DKA dataset and Wikipedia sample), we observe that the results on the DKA dataset are overall better, showing that the LLM is able to perform the task of omitting inferable terms very effectively, and lead to an easy reconstruction. Another interesting insight can be seen in the $\Theta$ metrics of the distilled text (i.e., matching distilled text with the ground thruth). In the Wikipedia sample evaluation, the $\Theta$-recall is 0 as we remove them directly from the text while in the DKA dataset evaluation the $\Theta$ recall is $8.6\%$, showing that the LLM does not erase all of the terms from the list, but only those it considers are easily inferable from the context. This result is also consistent with the exploratory analysis result shown in Table 1.
We do not see a significant difference among models, and any of them can be used effectively for this task. More closely, we observe how the T5 model obtains better results overall than Qwen2.5 in its 0.5B version, and are very comparable with its 1.5B. It beats the 1.5B version in some metrics even being nearly seven times smaller. For its smaller size, associated cost and overal good performance, it seems to be the sweet spot for this reconstruction model.
@@
$\Theta$ Precision (%) 91.50(6.98) 93.35(6.10) 88.28(8.27) 90.94(7.21) 96.85(14.92)
$\Theta$ Recall (%) 86.18(8.67) 87.14(8.60) 88.93(7.67) 92.04(6.53) 8.61(10.72)
$\Theta$ F1 (%) 88.55(6.92) 89.97(6.67) 88.44(7.13) 91.36(6.11) 14.25(15.76)
SacreBLEU (%) 85.09(8.10) 87.95(7.57) 86.76(7.80) 88.96(7.04) 49.16(11.59)
METEOR (%) 93.69(4.33) 94.65(4.00) 94.66(4.24) 95.59(3.62) 75.42(6.85)
ROUGE-1 (%) 95.99(2.51) 96.66(2.31) 96.00(2.81) 96.64(2.44) 85.86(4.10)
ROUGE-L(%) 94.95(3.18) 95.85(2.88) 94.95(3.47) 95.94(2.94) 85.86(4.10)
Cosine Similarity (%) 99.99(0.01) 100.00(00.01) 99.99(00.02) 99.99(00.02) 99.92(00.04)
Perplexity 19.35(6.83) 17.04(6.10) 18.42(6.68) 17.59(6.23) 99.77(61.36)
Perplexity Original 14.58(4.83) 14.58(4.83) 14.58(4.83) 14.58(4.83) 14.58(4.83)
Saved Tokens (%) 20.58(5.60) 20.58(5.60) 20.58(5.60) 20.58(5.60)
@@
Table 3: Average (standard deviation) metrics results over the Distilled Knowledge Answers dataset (DKA).
@@ 8.1 Cost Saving
Estimating the cost of the use of LLM API services at token level is very difficult GPT . GPT-x costs depends on many factors, including type of task, prompt quality, customization, API call volume, and computational resources. Nevertheless, the model size is the most relevant feature when considering model inference cost. Therefore, we want to make an exercise to illustrate how we should think about the potential cost savings of our pipeline. Our proposal should consider the gain/losses of processing text, which for cost purposes is measured by its size. The relevant text fragments are:
Extra inputs ($I$) : the cost of extending the prompt to specify the task of omitting function words. Specifically for our pipeline, the difference between size($Q$) and size($Q,\mathcal{P}(S)$).
Gained output ($G$): the gains from saved tokens compared to the vanilla generation. This is the difference between size($A$) and size($A_{D}$).
Reconstruction inputs ($I_{R}$): the cost of querying for the reconstruction task. This is equivalent to size($A_{D}$).
Reconstruction output ($O_{R}$): the cost of generating the final output. This is equivalent to size($A$).
@@
Assuming the costs of processing text for the reconstruction models, are two orders of magnitude cheaper, as seen in the evaluation, the relevant term in the right hand side is $C^{i}_{D}I$. Using available prices for the GPT-4o API service in its "gtp-4o-2024-08-06" version (i.e., \$2.5 per million input tokens, \$10 per million output tokens) as a proxy of cost, we see that the ratio $C^{i}_{D}/C^{o}_{D}$ is $\sim 0.25$. Therefore, any unrestricted domain knowledge answer will compensate processing $I$, which in our experiments corresponds to 97 extra words (See prompt 2 in Supplementary Material A.1).
@@ 8.2 Other language opportunities
Even though this study only focused on English, the technique is language independent. Several studies indicate that most languages also follow the Zipf distribution Piantadosi (2014), and that frequency ranks have a high correlation among languages Calude and Pagel (2011). However, these analysis were interested in word semantics rather than in function words. Moreover, our final inferable set $S$ only has four words in common with the first 100 words from the Swadesh vocabulary Swadesh (1952), a benchmark word list for language comparison. So, the particular benefits depending on the language is an open question, and we leave it as future work. We expect that languages such as Spanish, French, and Italian, which also use frequent function words to support grammatical cohesion, may adapt well to this methodology as they share structural similarities with English. However, challenges arise with languages that feature richer morphological complexity, like Finnish or Turkish, where grammatical meaning is embedded within word forms.
@@ 9 Conclusion
In this work we proposed an algorithm for the identification of a set of easily-inferable terms based on the context according to the desired cost-savings by using masking together with language models. We have proposed TRIM, a pipeline that allows cost savings in LLM-generation by omitting the generation of terms that do not contribute to the semantic meaning, and can be easily inferred a posteriori with a smaller language model. We have evaluated our approach in the question-answering knowledge domain, where we obtained a subset of functions words that are easily inferable. We have trained 4 different language models in the task of reconstructing a distilled answer. These models obtain very good performance across all metrics without noticeable losses, while saving 20.58% tokens with respect to the original full answer. We also showed that, while the output generation is more expensive than processing input prompts, in the vast majority of cases it will be better of to process knowledge answers through the cost-saving pipeline.
@@ 10 Limitations
Although the experiments explored the capability of an LLM to generate output without including terms that could be inferred later, it did not consider how this might negatively impact the LLM’s performance. LLMs are trained on vast datasets of human-written text that follow specific grammatical rules. Consequently, their output is based on this training data and on the sequence of tokens already produced during a session. If the usual grammatical structure is absent at generation and at the previous generated data, this could disrupt the ongoing token generation.
In a proper implementation, the potential penalty would only affect the generation of individual paragraphs in a response. After generating a paragraph, the reconstruction model would restore its grammatical structure. The distilled paragraph in the context window would then be replacer by the reconstructed one. For subsequent generations, as the intrinsic grammatical structure was restored, the LLM would not face any penalties regarding the previous generations.
Moreover, in this study we have primarily focused on the knowledge domain, where structured and factual content enables effective token-saving by omitting low-information words with minimal impact on semantic clarity. However, to generalize this methodology, it is necessary to explore its applicability in other domains with more variable language patterns, such as conversational dialogue where contextual dependencies and linguistic complexity could pose challenges. Testing across these diverse domains would provide a deeper understanding of the method’s adaptability and potential adjustments needed to preserve coherence and precision in different types of language processing tasks.
Finally, another relevant limitation is that we have used automated metrics such as SacreBLEU, ROUGE, and perplexity. This provide valuable insights into grammatical similarity, semantic coherence, and fluency, but they may not fully capture the nuances of human comprehension and acceptability Warstadt (2019), especially in reconstructed texts. A future human evaluation would be desired to assess subtle aspects like readability, naturalness, and the overall quality of the reconstructed responses, which metrics alone may overlook. Including human evaluators would enable a more comprehensive understanding of the method’s effectiveness, particularly in scenarios where slight changes in phrasing or word choice could impact clarity and user satisfaction.
@@
© 2024 JPMorgan Chase & Co. All rights reserved

### v5

@@ 7.2 Inferable Word Set
In this study, we obtain the set of words $W$ that will be used in the next experiment, analyzing how metrics degrade as the size of this set increases. This will allow us to empirically decide a good choice of $W$. We applied Algorithm 1 using a random sample of 100,000 paragraphs from the Wikipedia dump as the text corpus, an extended list of 127 English function words Jurafsky and Martin (2000) as word set, and RoBERTa as the language model. This returned the list of function words sorted by ease of inference based on the context regarding the text corpus.
@@
To determine the boundary for $W$, we followed an iterative procedure in which the T5 model is fine-tuned for the reconstruction task using the top-N words from the sorted list, increasing the subset in 5 words, from size 5 up to 60 (12 levels). For each of these trainings we used a new random Wikipedia sample containing 100,000 paragraph pairs (distilled, original) where the distilled text is built by removing the words of the corresponding level. Additional 5,000 pairs for testing were built with the same schema. The performance is then evaluated at each of these levels, analyzing the degradation of the results as the set size increases.
@@
Figure 2 shows the metric values plotted against the percentage of saved tokens per level, with trend lines extrapolated. From level 5 onward (about 18% saved tokens), further increases in saved tokens are minimal while metric degradation continues. Thus, we set the boundary at level 5 and select these 25 function words (set $W$) for the final evaluation.
@@ 7.3 Reconstruction Model
From this point forward, we use the fixed set $W$ and focus on developing the final reconstruction models $M_{R}$ to infer the omitted words. For training, we selected a new random sample of 100,000 Wikipedia paragraphs and removed the easily inferable words in $W$, creating the training corpus. The four chosen language models for $M_{R}$ were fine-tuned on this reconstruction task using the new corpus. Training details are provided in Table 7 of the Supplementary Material.
We tested each reconstruction model $M_{R}$ on two evaluation sets. First, on a new sample of 9,000 Wikipedia paragraphs to validate the training, and second, on the NaLDA dataset for each of the generation models $M_{G}$ (Section 6). Table 3 presents the NaLDA results with GPT-4o as $M_{G}$, including metrics comparing the distilled and ideal answers. Detailed results for other generation models are provided in Table 9 of the Supplementary Material.
Regarding the $\Theta$ metrics on the NaLDA dataset (Table 3), overall results indicate good performance (82.8 F1 score or higher), with no significant differences between models. Smaller encoder-decoder models like Bart and T5 show higher precision, while larger decoder-only models such as Qwen2.5 (0.5B and 1.5B) achieve better recall. These differences likely stem from their architectures: encoder-decoder models use cross-attention for better context understanding, while decoder-only models tend to be more verbose and infer more terms.
Regarding the rest of metrics, we observe that the grammatical metrics have good values across all models as well, being Qwen2.5(1.5B) the one obtaining the best results. All models achieve near-perfect scores on the semantic metric, demonstrating that their outputs effectively preserve the intended meaning. The perplexity suffers a small increase over the original one, being Qwen2.5(1.5B) the one obtaining the lowest perplexity. These results were obtained with a significant percentage of saved tokens, 18.85% in the Wikipedia sample and 19.41% for the answers in the NaLDA dataset.
@@
$\Theta$ Precision (%) 88.81(9.59) 90.14(9.41) 82.11(11.64) 85.35(10.78) 95.67(17.31)
$\Theta$ Recall (%) 82.50(11.04) 82.24(11.91) 84.33(10.27) 87.65(9.19) 8.61(10.50)
$\Theta$ F1 (%) 85.11(9.36) 85.51(9.79) 82.77(9.98) 86.10(9.12) 14.11(15.51)
SacreBLEU (%) 81.32(9.83) 82.92(10.69) 81.59(9.84) 83.99(9.25) 49.25(12.61)
METEOR (%) 91.51(5.79) 92.15(6.85) 92.73(5.42) 93.70(4.84) 74.95(7.52)
ROUGE-1 (%) 94.82(3.33) 95.13(3.99) 94.62(3.54) 95.26(3.25) 85.79(4.47)
ROUGE-L(%) 93.56(4.08) 93.96(4.66) 93.03(4.43) 94.10(3.98) 85.78(4.50)
Cosine Similarity (%) 99.98(0.03) 99.99(00.02) 99.98(00.02) 99.98(00.03) 99.92(00.04)
Perplexity 35.06(18.72) 31.53(17.46) 32.52(17.48) 31.13(16.54) 158.66(139.41)
Perplexity Original 25.59(13.87) 25.59(13.87) 25.59(13.87) 25.59(13.87) 25.59(13.87)
Saved Tokens (%) 19.41(5.84) 19.41(5.84) 19.41(5.84) 19.41(5.84)
@@
Table 3: Average (standard deviation) metrics results over the NaLDA dataset using GPT-4o as $M_{G}$
The metrics for the distilled text show significantly lower grammatical scores and higher perplexity, while the semantic metric remains nearly perfect. This indicates that removing the easily-inferable word set $W$ disrupts grammatical structure and coherence but largely preserves the original semantic content.
Since there is no significant difference among models, any can be used effectively for the reconstruction task. Notably, T5 outperforms Qwen2.5 (0.5B) and is comparable to Qwen2.5 (1.5B), even surpassing the latter in some metrics despite being nearly seven times smaller. Given its smaller size, lower cost, and strong performance, T5 appears to be the optimal choice for this task. Focusing on T5, we compared generation models $M_{G}$ for generalization (Table 4). All models perform the task successfully with similar results, despite large differences in price and size. However, comparing models is complicated due to variations in saved tokens, which are inversely related to metric performance. This highlights the value of prior experimentation (Section 7.2) in identifying the balance between saved tokens and desired performance, as some models are more verbose and use more function words. Overall, GPT-4o follows specific instructions most effectively, achieving the highest $\theta$ metrics, while Claude models show stronger grammatical metrics but save the fewest tokens.
GPT-4o GPT-4.1-mini GPT-4.1 Sonnet-4 Opus-4.1
$\Theta$ Precision (%) 90.14(9.41) 87.14(11.28) 88.41(10.10) 87.15(12.42) 88.56(9.98)
$\Theta$ Recall (%) 82.24(11.91) 79.64(12.39) 82.22(12.01) 81.56(13.74) 81.12(12.12)
$\Theta$ F1 (%) 85.51(9.79) 82.52(10.87) 84.73(10.06) 83.39(12.51) 84.03(10.32)
SacreBLEU (%) 82.92(10.69) 75.21(10.33) 81.78(9.48) 85.13(9.00) 86.49(8.25)
METEOR (%) 92.15(6.85) 88.49(6.01) 92.10(5.00) 93.48(5.26) 93.76(5.17)
ROUGE-1 (%) 95.13(3.99) 92.58(3.68) 94.86(3.09) 95.71(3.28) 96.11(3.09)
ROUGE-L (%) 93.96(4.66) 90.95(4.35) 93.75(3.70) 94.81(3.74) 95.22(3.50)
Cosine Similarity (%) 99.99(0.02) 99.98(0.02) 99.99(0.02) 99.99(0.02) 99.99(0.02)
Perplexity 31.53(17.46) 37.67(20.43) 35.03(18.20) 35.36(19.04) 35.25(18.36)
Perplexity Original 25.59(13.87) 26.51(16.13) 27.18(12.89) 30.85(18.06) 30.74(15.90)
Saved Tokens (%) 19.41(5.84) 22.79(5.44) 19.45(5.32) 17.83(5.00) 17.38(4.36)
Table 4: Average (standard deviation) metrics results over the NaLDA dataset using T5 as $M_{R}$
We also compared evaluation results between the NaLDA dataset (using GPT-4o as $M_{G}$) and a Wikipedia sample. Again, the results are consistent with LLM showing the ability to omit only contextually inferable words, rather than all pre-defined words, highlighting its nuanced understanding. See results for Wikipedia sample and discussion in Appendix A.5
7.3.1 Results across different task types
Examining results across the five NaLDA task types using GPT-4o as $M_{G}$ and T5 as $M_{R}$ (Figure 3), several insights emerge. The reconstruction model $M_{R}$ performs best on the knowledge-answer task, where the LLM generates responses to knowledge questions. These outputs are argumentative and align well with the Wikipedia-based training data. In contrast, $M_{R}$ shows lower performance on rewrite, summarize, and translate tasks, with an average metric decrease of about 5.77%. This suggests that the TRIM pipeline is more effective when the LLM produces free-form text rather than working from a provided passage.
Additionally, the percentage of saved tokens varies across task types, even though the word set $W$ remains constant. This variation does not directly correlate with evaluation metric performance. For instance, the knowledge-answer task type achieves both high scores and a high percentage of saved tokens, while the create task type saves fewer tokens but still performs well. This suggests that factors such as task type and the nature of the output text also influence token savings. Aside from these differences, all F1 scores are above 8̃0%, indicating that the pipeline is effective across all narrative tasks.
Figure 3: Comparison of metrics for GPT-4o as $M_{G}$ and T5 as $M_{R}$ across the 5 different task types of the NaLDA dataset.
@@ 8.1 Cost Saving
Estimating the cost of using LLM API services at token level is a non trivial task GPT . Costs depend on many factors, including type of task, prompt quality, customization, API call volume, and computational resources. Nevertheless, the model size is the most relevant feature when considering model inference cost. Neither OpenAI nor Anthropic have officially disclosed the size of their most recent models. There have been attempts of estimate the size of these models by comparing them with open-source models, reaching a general consensus estimate for the GPT-4o and Sonnet 3.5 models of around 200 billions parameters, placing models such as Sonnet 4, GPT-4.1 and Opus 4.1 at even higher values Abacha et al. (2025); Erdil (2024). Although not knowing their sizes exactly, we want to make an exercise to illustrate how we should think about the potential cost savings of our pipeline. Our proposal should consider the gain/losses of processing text, which for cost purposes is measured by its size. The relevant text fragments are:
Extra inputs ($I$) : The extra text added to the prompt to specify the task of omitting function words. Specifically for our pipeline, the size of $\mathcal{P}(I,W)$.
Gained output ($G$): the gains of saved tokens compared to the vanilla generation. This is the difference between size($A$) and size($A_{D}$).
Reconstruction inputs ($I_{R}$): the query for the reconstruction task. This is equivalent to size($A_{D}$).
Reconstruction output ($O_{R}$): the text generated for the final output. This is equivalent to size($A$).
@@
In our evaluation, reconstruction models range from 0.139 to 1.5 billion parameters, which is at least two orders of magnitude smaller than the size estimation of the generation models $M_{G}$ employed like GPT-4o, GPT-4.1, Sonnet 4 and Opus 4.1. Therefore, the relevant term in the right hand side is $C^{i}_{D}I$. Using available prices for the GPT and Claude API services as a proxy of cost, we see that the ratio $C^{i}_{D}/C^{o}_{D}$ is $\sim$0.25 for GPT and $\sim$0.2 for Claude. In our experiments, instruction $I$ adds 169 extra tokens to the prompt (See prompt 2 in Supplementary Material A.1), so it would be compensated after 42 and 34 omitted tokens (fewer number of words) for the GPT and Claude models respectively. The final results depends on the tokens saved by each model using TRIM, but to exemplify with GPT-4o (19.4% saved tokens) and Sonnet 4 (17.83% saved tokens), natural language generation answers longer than $\sim$216 tokens and $\sim$190 tokens will have better inference cost with our pipeline respectively.
@@ 8.2 Other Language Opportunities
Although this study focused on English, the technique is language independent. Most languages follow the Zipf distribution Piantadosi (2014), and frequency ranks are highly correlated across languages Calude and Pagel (2011), though prior analyses emphasized word semantics over function words. Our final easily-inferable word set $W$ overlaps with only four words from the Swadesh list Swadesh (1952), leaving language-specific benefits as an open question for future work. Languages like Spanish, French, and Italian, which rely on frequent function words for grammatical cohesion, may adapt well to this approach, while languages with greater morphological complexity, such as Finnish or Turkish, may present additional challenges.
@@ 9 Conclusion
We proposed TRIM, a pipeline for reducing LLM text generation costs by omitting words that are easily inferred and reconstructing them with a smaller model trained for this task. To support this, we introduced an algorithm for ranking word inferability, built the NaLDA evaluation dataset across five generation tasks, and presented a new reconstruction metric. Our experiments with four reconstruction models showed strong performance and an average token savings of 19.41% for GPT-4o, demonstrating TRIM’s potential for improving cost efficiency in natural language generation.
@@ 10 Limitations
We acknowledge some limitation this pipeline could have both during evaluation and in possible application scenarios out of scope. In the evaluation, we have attempted to cover a wide range of perspectives, from semantic content to grammatical structure evaluations, including the coherence of the text and a new metric specifically focused on the task of text reconstruction, paying attention to the subset of terms expected to be omitted and reconstructed. Thus, there are scenarios where automated evaluation may error and excessively penalize the effectiveness of the entire pipeline. Examples of this occurrence can happen in scenarios where function words in the text can be omitted or exchange with others while the sentence remains grammatically and semantically correct, preserving also its coherence. An example can be the sentence "I work from 9 to 5." If we compare it with the sentence "I work 9 till 5," the function word "from" has been omitted and the function word "to" has been replaced by "till." While the evaluation of semantic meaning will indicate that these two sentences are equivalent, the rest of the general and specific grammatical structure metrics will indicate lower performance. Therefore, a human evaluation of the entire pipeline would be interesting to assess these types of scenarios as well.
@@
Although the experiments explored the capability of an LLM to generate output without including words that could be inferred later, it did not consider how this might negatively impact the LLM’s performance. LLMs are trained on vast datasets of human-written text that follow grammatical rules. Consequently, their output is based on this training data and on the sequence of tokens already produced during a session. If the usual grammatical structure is absent at generation and at the previous generated data, this could disrupt the ongoing token generation. For example, a previous study remarks the importance of a (different) set of tokens at generation time called stylistic tokens (e.g., ‘Hello’, ‘Thank’, ‘However’, ‘Remember’, etc.), which include transitional phrases, discourse markers, and safety disclaimers Lin et al. (2023). Nevertheless, in a proper implementation, the potential penalty would only affect new generation of text. After generating a piece of text, the reconstruction model would restore its grammatical structure. The distilled piece of text in the context window would then be replaced by the reconstructed one. For subsequent generations, as the intrinsic grammatical structure was restored, the LLM would not face any penalties regarding previous generations.
In this study we have primarily focused on the natural language generation domain, where structured and verbose content enables effective token-saving by omitting low-information words with minimal impact on semantic clarity. However, generalizing this pipeline to other domains is not straightforward, like conversational as it could lack of enough amount of easily-inferable low-semantic words for the trade-off of adding extra tokens at the initial instruction, or code generation where one bad inferred word could imply an execution error. Moreover, during experimentation neither the so called "reasoning" models like o3 nor the reasoning capabilities of models like Clause Opus 4.1 or Sonnet 4 were used. Despite believing that the reasoning process can greatly benefit from the use of this pipeline from the cost perspective, the focus of this research has been on the response generated by the model itself and its applicability for the reasoning step is left for future research.
@@
© 2025 JPMorgan Chase & Co. All rights reserved